OpenStudy (anonymous):

The numbers 20, 21, and c are a Pythagorean triple. What is the value of c? A. 25 B. 41 C. 22 D. 29

OpenStudy (anonymous):

Whichever of these produces a c as an integer will be your answer: \(\large 20^2+21^2=c^2 \) or \(\large 20^2+c^2=21^2 \)

OpenStudy (anonymous):

@ByteMe 20^2 + 21^2 = 29^2?

OpenStudy (anonymous):

yep!!!

OpenStudy (anonymous):

thanks!